A cylindrical water tank carries a volume of 150 ℓ when it is full. The water tank has a base surface

with a diameter of 140 cm. What is the height of the water tank? Round your answer to three decimal

digits.

[1] 9,744 cm

[2] 0,097 cm

[3] 2,440 cm

[4] 0,024 cm


1
Expert's answer
2021-06-09T17:02:58-0400

"V = \pi r^2 h= \pi (\frac{d}{2})^2 h" so "h = \frac{4 V}{\pi d^2}",


150 L = 150 000 cm3 and "h = \frac{4 * 150000}{3.1416 *140^2} \approx 9.744 cm" 9.744